Emily and Andy each go to a hardware store to buy wire. The table shows the cost y in dollars for x inches of the wire they need. Emily needs 22 feet of the wire. Andy needs 13 yards of the wire. How much will each of them spend for​ wire?


1
Expert's answer
2022-02-03T16:36:26-0500

The given table is;


Length in Inches (x) | Cost in Dollars (y)

100 | 5.00

145 | 7.25

150 | 7.50

185 | 9.25


From the table above, 1inch costs $0.05 since;


"\displaystyle
\frac{5}{100}=\frac{7.25}{145}=\frac{7.50}{150}=\frac{9.25}{185}"


Now,

For Emily: Since"\displaystyle
1 \text{inch}=\frac{1}{12} \text{ foot, thus Emily needs }264\text{ inches of the wire.}\\"

But from the question, 1 inch of wire is sold at $0.05, thus we have that Emily will spend "\displaystyle
\$\left(\frac{264\times5}{100}\right)=\$13.2" for her wire.


For Andy: Since"\displaystyle
1 \text{inch}=\frac{1}{36} \text{ yard, thus Andy needs }468\text{ inches of the wire.}\\"

But from the question, 1 inch of wire is sold at $0.05, thus we have that Andy will spend "\displaystyle
\$\left(\frac{468\times 5}{100}\right)=\$23.40" for her wire.